2x + y = 6
3x + y = 12

Answer by rahman(247)

2x + y = 6
3x + y = 12
------------ -
-x=-6
x=6
Substitute x=6 to 2x + y = 6
2x + y = 6
2(6) + y = 6
12 + y = 6
y = 6 - 12
y = -6
Thus, x=6 and y=-6.